OpenStudy (anonymous):

x+y+x=53 3x-2y+z=69 -x+2y-z=-59 i got (68, -24, 9)

OpenStudy (anonymous):

Hmm... I actually got different answers. I might be wrong though. You want to work on this together?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i did y

OpenStudy (anonymous):

Oh wait...

OpenStudy (anonymous):

x+y+x=53 <--- are you sure it's two x's ? 3x-2y+z=69 -x+2y-z=-59

OpenStudy (anonymous):

last one a z my bad

OpenStudy (anonymous):

Ah ok. This will change up my answer a little. Regardless, here's what we shall do: equation 1 = x+y+z=53 We want to solve what Y is first, so... let's put y by itself. y = -x - z + 53

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Mmk. 3x-2y+z=69 3x = 2y - z + 69 x = (2y - z + 69)/3

OpenStudy (anonymous):

To be honest, though... z would be A LOT easier to use. Should we do that instead?

OpenStudy (anonymous):

Your choice.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Awesome. Equation 2: 3x-2y+z=69 z = -3x + 2y + 69 So far so good?

OpenStudy (anonymous):

yea writing the steps down

OpenStudy (anonymous):

Okie dokie. I'll go ahead and begin explaining the next step. We're going to combine this last equation into the equation we made earlier. y = -x - z + 53 This was equation 1 in the form of y = blahblah z = -3x + 2y + 69 This was equation 2 in the form of z = blahblah We're going to create what I call equation 1.2 (meaning equation 1 combined with 2) y = -x - (-3x + 2y + 69) + 53 Let me know if this step made sense. It's important that this makes sense before we do anything else.

OpenStudy (anonymous):

ok i dont get this step

OpenStudy (anonymous):

Okie dokie. Thanks for letting me know. ^_^ Remember how equation one was written to say y = -x - z + 53 , right?

OpenStudy (anonymous):

And how we made equation 2 say z = -3x + 2y + 69 ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Alright. Because equation two says, "the letter z is exactly "-3x + 2y + 69", we can replace any z that we see with "-3x + 2y + 69". Does that statement make sense?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Snazzy. So what we did is go back to equation 1: y = -x - z + 53 And replaced the z with what?

OpenStudy (anonymous):

-3x + 2y + 69

OpenStudy (anonymous):

Excellent. So we're left with y = -x - (-3x + 2y + 69) + 53 right?

OpenStudy (anonymous):

(The stuff in parentheses replacing the original z that was there)

OpenStudy (anonymous):

yea i see

OpenStudy (anonymous):

Alright. Now we're going to distribute that minus that's in front of our parentheses.

OpenStudy (anonymous):

y = -x - (-3x + 2y + 69) + 53 y = -x +3x -2y - 69 + 53 Looking good so far?

OpenStudy (anonymous):

yea and sorry idk why it keeps putting i dont get this step every time i post

OpenStudy (anonymous):

Yeah, I noticed lol. It's ok, this site does that kind of stuff. :P

OpenStudy (anonymous):

Now we combine similar terms: y = -x +3x -2y - 69 + 53 y = 2x -2y - 16 (be careful that you don't accidentally say you have 4x) 3y = 2x - 16 So far so good?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Awesome. Now, we're going to want to note this equation, because it's missing a Z, which is great. So, let's call it equation 1.2 (as in equation 1 mixed with equation 2) ***** Equation 1.2 ******* 3y = 2x - 16 I'll show you in a few moments why we are going to like this equation.

OpenStudy (anonymous):

Alright. Last thing we want to do is get x by itself in equation 3. Equation 3: -x+2y-z=-59

OpenStudy (anonymous):

-x+2y-z=-59 2y -z + 59 = x x = 2y - z + 59 Makes sense so far?

OpenStudy (anonymous):

yea and thanks for going slow I starting to get it

OpenStudy (anonymous):

Not a problem! I'm just happy that you're willing to learn instead of just demanding the answer. :D

OpenStudy (anonymous):

Alright, so remember in our awesome equation 1.2, we only had x and y?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

3y = 2x - 16 What would happen if we plugged in x = 2y - z + 59 You should notice a problem that would occur.

OpenStudy (anonymous):

As in, 3y = 2x - 16 3y = 2(2y - z + 59) - 16

OpenStudy (anonymous):

we can simplify the y?

OpenStudy (anonymous):

We technically can, but there's a bigger problem. I'll give you a hint: "Alright, so remember in our awesome equation 1.2, we only had x and y?" Why do you think I asked that question? :P

OpenStudy (anonymous):

In other words, we started with 3y = 2x - 16 where there were only two variables. But once we put in our x = 2y - z + 59, we have a problem.

OpenStudy (anonymous):

z came into the equation?

OpenStudy (anonymous):

That's right. How do you think we will make z go away and replace it with something nicer?

OpenStudy (anonymous):

ok i dont get this step

OpenStudy (anonymous):

I would say cancel it out but I don't know how that would work

OpenStudy (anonymous):

Well, we did something earlier that let us get rid of the original z. Do you recall what we did?

OpenStudy (anonymous):

Hint: Scroll up about halfway to where I said, "Snazzy. So what we did is go back to equation 1: y = -x - z + 53 And replaced the z with what?" And you answered it correctly.

OpenStudy (anonymous):

oh we replaced z with -3x + 2y + 69

OpenStudy (anonymous):

Exactly. :D So here's what we have now, after switching out x: 3y = 2(2y - z + 59) - 16 What will our new problem look like after we replace z?

OpenStudy (anonymous):

Hint: Make sure that you put the stuff that's replacing z in parentheses, or your negative won't carry through correctly.

OpenStudy (anonymous):

um 3y=2(2y-(-3x=2y+69)+59)-16?

OpenStudy (anonymous):

3y = 2(2y - (-3x + 2y + 69) + 59) - 16

OpenStudy (anonymous):

Though, I'm guessing your keyboard messed up on that = instead of +, right?

OpenStudy (anonymous):

yea i think

OpenStudy (anonymous):

Awesome. We're almost done! Now we just begin simplifying. 3y = 2(2y - (-3x + 2y + 69) + 59) - 16 3y = 2(2y +3x - 2y - 69 + 59) - 16 So far so good? Please double check this to be safe. It'd suck if we get messed up on this step.

OpenStudy (anonymous):

it looks good when i went over it i didnt see anything that looked wrong

OpenStudy (anonymous):

Awesome. 3y = 2(2y +3x - 2y - 69 + 59) - 16 3y = 2(0y + 3x - 10) - 16 3y = 0y + 6x -20 - 16 3y = 6x -36 So far so good?

OpenStudy (anonymous):

Note that the 0y can just be removed since "no y's" is the same as "nothing".

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Awesome. Now, notice there's something we can do on this step to make it a lot prettier. What do you think that is?

OpenStudy (anonymous):

(Think greatest common factors)

OpenStudy (anonymous):

3y = 6x -36

OpenStudy (anonymous):

something with 6 and 36?

OpenStudy (anonymous):

Something even easier. Hint: It has to do with a 3.

OpenStudy (anonymous):

3 and 6?

OpenStudy (anonymous):

We can just divide everything by 3, right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

3y = 6x -36 y = 2x - 12 Looks a lot nicer, ya?

OpenStudy (anonymous):

yea it does

OpenStudy (anonymous):

Alright, here's where it gets a little challenging. I'm going to let you figure this out (I'll give hints if you need them). We want to replace x with something that has just constants and a y, so that we can have only one variable. You know that we have to have an equation in the form of "x = yblahblah". Which equation will we use to get such a thing? I'll give you a hint if you need it.

OpenStudy (anonymous):

In other words, besides the one we're working with right now, which equation has only x's and y's? We do NOT want any z's messing with us.

OpenStudy (anonymous):

3y = 6x -36?

OpenStudy (anonymous):

But that's the same equation we're working with. If we put that in there, we'd get something that will lead us to 0=0 because everything would cancel out. We need a different equation.

OpenStudy (anonymous):

Hint: I said something about a "super awesome equation" earlier, right? ;) Now's her time to show us her magic!

OpenStudy (anonymous):

I didnt see it. Gimmy min let me look through my notebook.

OpenStudy (anonymous):

I'll go ahead and give it to you. 3y = 2x - 16 This was that special 'equation 1.2' that had only two variables. Remember that one?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Ok. So what are we going to do to that equation (3y = 2x - 16) to help us solve (y = 2x - 12) ?

OpenStudy (anonymous):

plug y in?

OpenStudy (anonymous):

Here's a useful hint: NORMALLY, you want to get it in the form of x = blah blah. But in this case, you can find a similar variable in both equations. Do you see anything repeated exactly the same?

OpenStudy (anonymous):

Hmm.... actually, we could do that too. I'll show you both ways.

OpenStudy (anonymous):

Method 1: (what I was planning on doing) Since 2x appears in both equations, you could do this: given: (3y = 2x - 16) and (y = 2x - 12) 3y = 2x - 16 3y + 16 = 2x 2x = 3y + 16 Since 2x = 3y + 16, we can change equation 1.2 from (y = 2x - 12) into (y = 3y + 16 - 12) Then simplify into (y = 3y + 4) (-2y = 4) y = -2 Does that make sense?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Method 2: (your easier way :P) given: 3y = 2x - 16 and y = 2x - 12 Plug in y into the first equation: 3(2x - 12 ) = 2x - 16 6x - 36 = 2x - 16 4x = 20 x = 5

OpenStudy (anonymous):

Oooo nice! Do you see what happened?

OpenStudy (anonymous):

you found what x = so no you can plug it in or something else?

OpenStudy (anonymous):

Yup! Let's use the easiest equation for that. x+y+x=53

OpenStudy (anonymous):

Oops. I copied it from the original question. Here's the correct one: x+y+z=53

OpenStudy (anonymous):

so be 5+y+z=53?

OpenStudy (anonymous):

But we also found y earlier, remember?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

5+-2+z=53

OpenStudy (anonymous):

Perfect. So 3 + z = 53. Now put in the last jigsaw on this beautiful puzzle. :P

OpenStudy (anonymous):

z=50 but thats not to big?